If n=26, ¯xx¯(x-bar)=31, and s=5, find the margin of error at a 99% confidence level
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 31

sample standard deviation = s = 5

sample size = n = 26

Degrees of freedom = df = n - 1 = 26 - 1 = 25

At 99% confidence level

= 1 - 99%

=1 - 0.99 =0.01

/2 = 0.005

t/2,df = t0.005,25 = 2.787

Margin of error = E = t/2,df * (s /n)

= 2.787 * ( 5 / 26)

Margin of error = E = 2.73